OpenStudy (casadie1004):

Does anyone know how to tell how many solutions something has? The equation is 2x=-10y+6, x=5y=3 Im not the best at math so if you could explain it in steps that would be amazing! fan and medal.

OpenStudy (danjs):

both of those are equations for a line the solution to a system of two equations in x and y, is the value or point (x,y) where the lines cross, there they will both have the same x and y, and both equations work

OpenStudy (danjs):

lines can have 1,0,or infinity solutions 1 if they cross each other 0 if they are parallel infinity if you have the same line written twice as a multiple of the other equation

OpenStudy (danjs):

overlapping lines intersect infinity times i guess

OpenStudy (casadie1004):

so how do i figure out the answer?

OpenStudy (danjs):

can you type those 2 equations again

OpenStudy (casadie1004):

2x=-10y=6, x+5y=3

OpenStudy (danjs):

the second has two = , not sure if that is right

OpenStudy (casadie1004):

Its not, i didnt realize my shift key got jammed.

OpenStudy (danjs):

so it is 2x = -10y + 6 x + 5 y = 3 there are a number of ways of figuring this out, you can graph them and see what it looks like without solving, a number of ways of solving , do you remember which way you have done maybe before?

OpenStudy (casadie1004):

i usually solved it, i tried graphing it once and was stuck, But i cant remember how to solve it. Its been awhile.

OpenStudy (danjs):

an easy way for 2 equations is to solve both of them for the same thing, x or y ill solve for x on both x = -5y + 3 x = -5y - 3

OpenStudy (danjs):

now you have x=both those things, so you can set them equal x = -5y + 3 x = -5y - 3 so -5y + 3 = -5y - 3 , add 5y to both sides 3 = -3 false statement, meaning no solutions to this one, the lines are parallel and never intersect

OpenStudy (danjs):

when you solve and get a false equality like that 3=-3, no solutions, parallel lines if you get a true statement like 5=5, then infinite solutions, both are the same lines overlapping

OpenStudy (danjs):

otherwise , you will get a unique x and y, the point of intersection

OpenStudy (casadie1004):

so do u not have to do the other equation now?